# Automatic Reviewers Fail to Detect Faulty Reasoning in Research Papers: A New Counterfactual Evaluation Framework

### v1

@@ Abstract
Large Language Models (LLMs) have great potential to accelerate and support scholarly peer review and are increasingly used as fully automatic review generators (ARGs). However, potential biases and systematic errors may pose significant risks to scientific integrity; understanding the specific capabilities and limitations of state-of-the-art ARGs is essential. We focus on a core reviewing skill that underpins high-quality peer review: detecting faulty research logic. This involves evaluating the internal consistency between a paper’s results, interpretations, and claims. We present a fully automated counterfactual evaluation framework that isolates and tests this skill under controlled conditions. Testing a range of ARG approaches, we find that, contrary to expectation, flaws in research logic have no significant effect on their output reviews. Based on our findings, we derive three actionable recommendations for future work and release our counterfactual dataset and evaluation framework publicly. [^1] https://github.com/UKPLab/counter-review-logic
@@ 1 Introduction
  • A model of paper soundness and research logic
@@ 3 Research Logic
As in any empirical science, a sound NLP or AI paper makes a contribution to collective knowledge and justifies it through valid reasoning based on empirical evidence Armstrong and Green (2022). By this definition, a paper’s soundness comprises four building blocks, which we refer to throughout the rest of the paper: the method describing the experimental setup, the empirical results generated by them, the conclusions drawn from these results, and the finding summarizing the conclusions as a scientific contribution. Well-designed papers explicitly present all four components. While not necessarily all papers present empirical findings, e.g., those with theoretical proofs or opinions, our analysis in Section˜5 shows that a vast majority of NLP and AI research papers make at least some empirical contribution. Formally, a paper $p$ claims to contribute findings $F$. The paper $p$ is sound only if all findings $f\in F$ are sound. A finding $f$ is sound iff all underlying conclusions $C(f)$ are sound and jointly support $f$. A conclusion $c\in C(f)$ is sound iff its underlying results $R(c)$ are sound and jointly support $c$. A result is sound iff the experimental methodology $m$ adheres to best scientific practice and plausibly produces the result. We assume that $C(f)$ and $R(c)$ are the minimal sufficient sets of conclusions and results, respectively; i.e., all elements are necessary to support the corresponding finding $f$ or conclusion $c$. These building blocks form a hierarchical, logical structure, which we call the paper’s research logic, in which arcs represent the support relation. Figure 1 illustrates this structure. Importantly, the research logic differs from the paper’s argumentation structure, which consists of rhetorical moves to make the paper appear interesting and plausible Teufel et al. (2009). In other words, soundness and convincingness are independent concepts of a paper.
Verification
@@ 4 Framework
We use ARG as an umbrella term for systems that receive a research paper as an input and output a review report. While ARGs usually involve LLMs, we do not probe LLM per-se but focus on their behavior in the context of automatic reviewing. A reliable ARG should detect flaws in research logic and reflect them in its output. Unlike prior work (Zhang and Abernethy, 2025, i.a.), which checks for mentions of specific issues, we focus on the average effect of flawed logic on automated reviews, recognizing that peer reviews weigh multiple factors and cannot highlight every concern. For instance, limited novelty might overshadow methodological flaws. Formally, we estimate the concept average treatment effect (ATE) Goyal et al. (2019) of research logic interventions on automatically generated reviews using approximate counterfactuals Gat et al. (2024). In other words, we surgically edit high-quality papers to compromise their research logic and quantify the resulting changes in automatic reviews, which should, on average, become more critical and emphasize soundness more compared to the automatic review of the original paper.
@@ 4.1 Pipeline
Our framework consists of three stages (Figure 2). First, given a paper $p$, we generate two sets of counterfactuals: soundness-critical $\text{CF}_{\text{CR}}(p)=\{p^{\prime}_{1},...,p^{\prime}_{m}\}$, which introduces errors to the paper’s research logic, altering its soundness while preserving other concepts such as clarity and novelty; and soundness-neutral $\text{CF}_{\text{NE}}=\{p_{1}^{*},...,p_{k}^{*}\}$, which apply surface-level edits (e.g., formatting and language changes) to later contextualize the effects of the soundness-critical edits.
@@ 4.2 Counterfactual Paper Generation
Table 1: Examples of perturbation operations targeting different building blocks from our dataset (shortened for brevity). We show the original building block, the compromised version, and an example paper edit.
@@ 5 Dataset
Applying our pipeline, we first create a dataset of papers, associated research logic, and counterfactual versions, and validate each step. We list the model versions for all steps in Appendix A.3 and prompts in the supplementary materials.
@@ 5.2 Research Logic Data
Through manual refinement, we develop the extraction prompts for the research logic on the development set and select the best performing model. GPT-4o-mini OpenAI et al. (2024) offered the best trade-off between inference speed, cost and output quality.
@@ 5.3 Counterfactual Data
We tune prompts and select the best LLM based on manual inspection on the development set.
@@
We generate soundness-neutral counterfactuals using Phi-4 14B Abdin et al. (2024). For active-to-passive and American-to-British, we randomly select 40% of the paragraphs; for language error, we select 20% to simulate minor language issues. Given the simplicity of these edits, one author manually inspects $10$ counterfactuals per type to verify that the edits do not affect the paper content. In all cases, the revised text preserves the original meaning.
@@
We use GPT-4o-mini to generate the soundness-critical counterfactuals and tune prompts on the development set. If an edit type is not applicable, e.g., a paper lacks the necessary claim types for finding edits, we skip the counterfactual for that paper.
@@ 5.4 Dataset Overview
Table 2: Dataset statistics for soundness-critical ($\text{CF}_{\text{CR}}$) and soundness-neutral ($\text{CF}_{\text{NE}}$) counterfactuals (CFs). The number of edits counts edit operations; diff. is the Levensthein distance.
Table 2 summarizes the final evaluation dataset comprising $931$ counterfactuals across $133$ papers, excluding $5$ papers without empirical findings. The number of edits and amount of changed text is comparable for both counterfactual types, ensuring fair comparison. With respect to diversity, among soundness-critical edits, $85.6\%$ of them appear in the text, $9.6\%$ in tables, and $4.7\%$ in figure captions; the latter two only occur for result edits. The dominance of text edits reflects the textual nature of findings and conclusions, and the fact that many results are reported and interpreted in the text. Each edit type engages distinct reasoning over different areas of the paper.
@@ 6.1 Experiments
We consider three types of ARGs from the literature. The first type uses LLMs in zero-shot mode either with a generic review prompt (Zero-Generic), e.g. in Liang et al. (2024b), or a prompt including venue-specific guidelines (Zero-Guide), e.g. in Du et al. (2024). We test both with GPT-4o-mini and GPT-4.1 as large proprietary LLMs, DeepSeek-14B and DeepSeekV3 Liu et al. (2024) as reasoning LLMs, and Phi-4 as a small open-weight model. The second type comprises multi-agent systems where LLMs specialize in different paper aspects and engage in a discussion. We include TreeReviewer Chang et al. (2025), implemented with GPT-4o-mini, as a representative. MARG D’Arcy et al. (2024), evaluated on our development set, proved computationally infeasible with individual reviews requiring at least $20$ minutes and up to $2$ hours. The third type uses LLMs fine-tuned on peer review data. We test Reviewer2 Gao et al. (2024), fine-tuned with an automatic prompting model, and DeepReviewer Zhu et al. (2025), which is further trained on synthetic reasoning data. We use default hyperparameters for all ARGs. We fix the random seed and, for zero-shot LLMs, set the temperature to zero to minimize output variance. We truncate the paper to the effective context window size per LLM Hsieh et al. (2024). When ARGs produce semi-structured text, e.g. DeepReviewer, we parse the output using regular expressions and, if needed, fall back to GPT-4o-mini for parsing. We include reviews that cannot be parsed into the venue’s template in raw form; if the parsing of scores fails, we do not consider those scores for evaluation.
@@ 6.2 Results
Figure 3: The ATE of review difference for soundness-critical and soundness-neutral counterfactuals by aspects, sentiment, and score. The larger the difference between neutral and critical means the better.
@@
Figure 3 is a box plot of the ATE for each difference dimension and both counterfactual types per ARG. To recap, ’Aspect’ measures the number of soundness-related statements, ’Sentiment’ the density of positive comments, and ’Score’ the overall rating; all of them compare the ARG’s reviews for the counterfactuals with the one for the original (see Section˜4.3). For an ideal ARG, the ATE for soundness-critical and -neutral edits should lie far apart on all dimensions; the ATE of critical edits should be positive for aspects (soundness-related aspects are added) and negative for sentiment (less positive comments) and score (lower rating).
@@
Although none of the ARGs are significantly affected by flawed research logic, we rank the models to identify tendencies. A coherent ARG should show low ATE variance for soundness-neutral counterfactuals, being resilient to surface-level changes, and a large difference between the ATE for neutral and critical counterfactuals. To capture this in a single score, we express the ATE difference between both conditions as a multiple of the effect standard deviation of the neutral counterfactuals. For each ARG, we compute $z=\frac{|\text{ATE}_{\text{CR}}-\text{ATE}_{\text{NE}}|}{\sigma(\text{ATE}_{\text{NE}})}$ per dimension and then average over all dimensions. This yields the ranking shown in Table˜3. Notably, generic zero-shot prompts perform better than those with guidance, which may act as a distractor. For Phi-4, with its smaller context window, Zero-Guide likely causes truncation of key paper content. While the fine-tuned Reviewer2 ranks highest, DeepReviewer ranks lowest; likely, their backbone models, Llama-2 for Reviewer2 and Phi-4 for DeepReviewer, play an important role since the Zero-Guide results for Phi-4 suggest issues with context size and distractors. This warrants further investigation in future work. Finally, there is no clear link between model size and performance, as Phi-4 and GPT-4.1 both rank highly.
@@
Notably, all ARGs show high standard deviations for soundness-neutral counterfactuals, suggesting that ARGs are sensitive to spurious features similar to issues of prompt sensitivity in LLMs Sclar et al. (2024). To examine this further, we measure lexical similarity using ROUGE-2 and use GPT-4o-mini to detect whether assertions align between the original and soundness-neutral reviews, computing the Jaccard index on assertions. We perform human validation of the automatic alignment with two annotators on $50$ assertion pairs with substantial agreement (Krippendorff’s $\alpha=0.74$) attesting an accuracy of $0.78$. We analyze $48$ randomly sampled papers (roughly 8 per conference). Table 5 in the Appendix reports the detailed results. Surface similarity ranges from $0.25$ ROUGE-2 for DeepReviewer to $0.54$ for Zero-Guide-GPT4om, indicating substantial changes in wording for many ARGs. This pattern holds for the assertion overlap, with a Jaccard index from $0.27$ for DeepReviewer to $0.57$ for Zero-Generic-GPT4om indicating low to moderate overlap in the contents of the reviews. Overall, GPT-4o-mini produces more similar reviews than other ARGs, regardless of the prompting strategy. Even allowing for some measurement error, ARGs remain highly sensitive to spurious surface changes in the input text. Our findings, point to another recommendation for future work: all ARG evaluations, whether using human reference data or sensitivity analysis, should report mean performance on multiple soundness-neutral versions of the papers akin to prompt-sensitivity reporting to better reflect their true capability independent of spurious features. Here, our counterfactual dataset may also help improve model consistency through training data augmentation Wu et al. (2021).

### v2

@@ Abstract
Large Language Models (LLMs) have great potential to accelerate and support scholarly peer review and are increasingly used as fully automatic review generators (ARGs). However, potential biases and systematic errors may pose significant risks to scientific integrity; understanding the specific capabilities and limitations of state-of-the-art ARGs is essential. We focus on a core reviewing skill that underpins high-quality peer review: detecting faulty research logic. This involves evaluating the internal consistency between a paper’s results, interpretations, and claims. We present a fully automated counterfactual evaluation framework that isolates and tests this skill under controlled conditions. Testing a range of ARG approaches, we find that, contrary to expectation, flaws in research logic have no significant effect on their output reviews. Based on our findings, we derive three actionable recommendations for future work and release our counterfactual dataset and evaluation framework publicly. [^1] https://github.com/UKPLab/tacl2026-counter-review-logic
@@ 1 Introduction
  • A formal model of paper soundness as a product of its underlying research design and reasoning, i.e., its research logic.
@@ 3 Research Logic
As in any empirical science, a sound NLP or AI paper makes a contribution to collective knowledge and justifies it through valid reasoning based on empirical evidence Armstrong and Green (2022). By this definition, a paper’s soundness comprises four building blocks, which we refer to throughout the rest of the paper: the method describing the experimental setup, the empirical results generated by them, the conclusions drawn from these results, and the finding summarizing the conclusions as a scientific contribution. Well-designed papers explicitly present all four components. While not necessarily all papers present empirical findings, e.g., those with theoretical proofs or opinions, our analysis in Section˜5 shows that a vast majority of NLP and AI research papers make at least some empirical contribution. Based on this, we define our model of soundness formally. Our model effectively operationalizes Bacon’s (1878) theory of the scientific method as inductive reasoning from evidence.
Formal Model
A paper $p$ claims to contribute findings $F$. The paper $p$ is sound only if all findings $f\in F$ are sound. A finding $f$ is sound iff all underlying conclusions $C(f)$ are sound and jointly support $f$. A conclusion $c\in C(f)$ is sound iff its underlying results $R(c)$ are sound and jointly support $c$. A result is sound iff the experimental methodology $m$ adheres to best scientific practice and plausibly produces the result. We assume that $C(f)$ and $R(c)$ are the minimal sufficient sets of conclusions and results, respectively; i.e., all elements are necessary to support the corresponding finding $f$ or conclusion $c$. These building blocks form a hierarchical, logical structure, which we call the paper’s research logic, in which arcs represent the support relation. Figure 1 illustrates this structure. Importantly, the research logic differs from the paper’s broader argumentative structure, which consists of rhetorical moves to make the paper appear interesting and plausible Teufel et al. (2009). In Aristotelian terms, the research logic refers to the logical appeal (logos) while the broader argumentation in the paper often also invokes ethos (credibility) and pathos (appeal to emotion). In other words, soundness and convincingness are independent concepts of a paper.
Soundness Verification
@@ 4 Framework
We use ARG as an umbrella term for systems that receive a research paper as an input and output a review report. While ARGs usually involve LLMs, we do not probe LLMs per-se but focus on their behavior in the context of automatic reviewing. A reliable ARG should detect flaws in research logic and reflect them in its output. Unlike prior work (Zhang and Abernethy, 2025, i.a.), which checks for mentions of specific issues, we focus on the average effect of flawed logic on automated reviews, recognizing that peer reviews weigh multiple factors and cannot highlight every concern. For instance, limited novelty might overshadow methodological flaws. Formally, we estimate the concept average treatment effect (ATE) Goyal et al. (2019) of research logic interventions on automatically generated reviews using approximate counterfactuals Gat et al. (2024). In other words, we surgically edit high-quality papers to compromise their research logic and quantify the resulting changes in automatic reviews, which should, on average, become more critical and emphasize soundness more compared to the automatic review of the original paper.
@@ 4.1 Pipeline
Our framework consists of three stages (Figure 2). First, given a paper $p$, we generate two sets of counterfactuals (CFs): soundness-critical $\text{CF}_{\text{CR}}(p)=\{p^{\prime}_{1},...,p^{\prime}_{m}\}$, which introduces errors to the paper’s research logic, altering its soundness while preserving other concepts such as clarity and novelty; and soundness-neutral $\text{CF}_{\text{NE}}=\{p_{1}^{*},...,p_{k}^{*}\}$, which apply surface-level edits (e.g., formatting and language changes) to later contextualize the effects of the soundness-critical edits.
@@ 4.2 Counterfactual Paper Generation
Table 1: Examples of perturbations targeting different building blocks from our dataset (shortened for brevity). We show the original building block, the compromised version, and resulting paper edit.
@@ 5 Dataset
Applying our pipeline, we first create a dataset of papers, associated research logic, and counterfactual versions, and validate each step. We list the model versions in Appendix A.3 and prompts in the supplementary materials. We develop prompts for each step individually and manually verify outputs on a subset of papers. We begin with a basic prompt and iteratively refine it using Claude Sonnet 3.5 Anthropic (2024) akin to meta-prompting Zhou et al. (2023). We adjust prompts until we consider the outputs correct for all test samples. All prompts follow a similar structure and use json output format in line with prompt engineering best practices Phoenix and Taylor (2024) (see Figure 6 in the Appendix).
@@ 5.2 Research Logic Data
Through manual refinement, we develop the prompts for extracting the research logic on the development set and select the best performing model. GPT-4o-mini OpenAI et al. (2024) offered the best trade-off between inference speed, cost and output quality. We employ self-refinement Madaan et al. (2023), which notably improved the alignment between building blocks consistent with findings on other reasoning-intensive tasks Sahoo et al. (2024). A single refinement cycle achieved the best performance–cost trade-off. We query the OpenAI API [^6] https://platform.openai.com/ for model inference. On average, the extraction of the research logic takes roughly five minutes per paper.
@@ 5.3 Counterfactual Data
Following our general prompt development paradigm, we tune prompts and select the best LLM based on manual inspection on the development set.
@@
We generate soundness-neutral counterfactuals using Phi-4 14B Abdin et al. (2024). We run Phi-4 (with Q4_K_M quantization) on two L40 GPUs with approximately 14GB effective memory use during inference. On average, the generation of one soundness-neutral counterfactual takes roughly three minutes. For active-to-passive and American-to-British, we randomly select 40% of the paragraphs; for language error, we select 20% to simulate minor language issues. Given the simplicity of these edits, one author manually inspects $10$ counterfactuals per type to verify that the edits do not affect the paper content. In all cases, the revised text preserves the original meaning.
@@
We use GPT-4o-mini to generate soundness-critical counterfactuals. During prompt development, we found that decomposing the process into multiple stages, where each corresponds to a single LLM call producing an intermediate result with an explanation, substantially improved output quality. We explicitly separate stages that require reasoning from those that modify the paper’s content since this produced more diverse outputs. We further divide creative stages, such as proposing new hypothetical results (see Sec. 4.2), into candidate generation and subsequent selection. If an edit type is not applicable, e.g., a paper lacks the necessary claim types for finding edits, we skip the counterfactual for that paper. The generation of one soundness-critical counterfactual takes on average roughly two minutes using the OpenAI API.
@@ 5.4 Dataset Overview
Paper Distribution
#papers $133$
#papers p. conference $22.50\pm 1.80$
#papers p. institution $1.38\pm.08$
Research Logic Distribution
#paper types 77
#papers p. paper type $19.29\pm 33.68$
#findings p. paper $3.69\pm 1.93$
Table 2: Dataset statistics for the papers and their extracted research logic in the test set.
@@
Table 3: Dataset statistics for soundness-critical ($\text{CF}_{\text{CR}}$) and soundness-neutral ($\text{CF}_{\text{NE}}$) counterfactuals. The number of edits counts edit operations; diff. is the Levensthein distance.
Table 2 and 3 summarize the final evaluation dataset of $931$ counterfactuals on $133$ papers, excluding $5$ papers without empirical findings.
Diversity of Counterfactuals
As shown in Table 3, the number of edits and amount of changed text is comparable for both counterfactual types, ensuring fair comparison. Among soundness-critical edits, $85.6\%$ of them appear in the text, $9.6\%$ in tables, and $4.7\%$ in figure captions; the latter two only occur for result edits. The dominance of text edits reflects the textual nature of findings and conclusions, and the fact that many results are reported and interpreted in the text. Each edit type engages distinct reasoning over different areas of the paper.
Diversity of Papers
As shown in Table 2, the papers are nearly evenly distributed across conferences. To assess authorship diversity, we use the last author’s affiliated institution as a proxy, retrieving metadata via the OpenAlex API [^7] https://openalex.org/ (excluding $10$ missed authors). On average, $1.38$ papers originate from the same institution, indicating high diversity. Regarding research logic, of the $7$ paper types classified during extraction, papers introducing new methodology dominate ($75\%$), followed by analysis ($12\%$) and dataset papers ($6\%$). This skew towards methodological work likely reflects the actual distribution of publications in AI and NLP conferences and encourages future work on more diverse paper types (see Sec. 7).
@@ 6.1 Experiments
We consider three types of ARGs from the literature. The first type uses LLMs in zero-shot mode either with a generic review prompt (Zero-Generic), e.g. in Liang et al. (2024b), or a prompt including venue-specific guidelines (Zero-Guide), e.g. in Du et al. (2024). We test both with GPT-4o-mini and GPT-4.1 as large proprietary LLMs, DeepSeek-14B and DeepSeekV3 DeepSeek-AI et al. (2024) as reasoning LLMs, and Phi-4 as a small open-weight model. We design the prompts based on related literature to ensure alignment, while adding a standardized output format for easy parsing. We perform formatting and plausibility checks on the development set without further tuning. The second type comprises multi-agent systems where LLMs specialize in different paper aspects and engage in a discussion. We include TreeReviewer Chang et al. (2025), implemented with GPT-4o-mini, as a representative. MARG D’Arcy et al. (2024), evaluated on our development set, proved computationally infeasible with individual reviews requiring at least $20$ minutes and up to $2$ hours. The third type uses LLMs fine-tuned on peer review data. We test Reviewer2 Gao et al. (2024), fine-tuned with an automatic prompting model, and DeepReviewer Zhu et al. (2025), which is further trained on synthetic reasoning data. We use default hyperparameters for all ARGs. We fix the random seed and, for zero-shot LLMs, set the temperature to zero to minimize output variance. We truncate the paper to the effective context window size per LLM Hsieh et al. (2024). When ARGs produce semi-structured text, e.g. DeepReviewer, we parse the output using regular expressions and, if needed, fall back to GPT-4o-mini for parsing. We include reviews that cannot be parsed into the venue’s template in raw form; if the parsing of scores fails, we do not consider those scores for evaluation.
@@ 6.2 Results
Figure 3: ATE of aspect differences for soundness-critical and -neutral CFs; the larger the difference of neutral/critical means (dashed horizontal lines) the better.
Figure 4: ATE of sentiment differences for soundness-critical and -neutral CFs; the larger the difference of neutral/critical means (dashed horizontal lines) the better.
Figure 5: ATE of score differences for soundness-critical and -neutral CFs; the larger the difference of neutral/critical means (dashed horizontal lines) the better.
@@
Figure 3, 4, and 5 show the box plots of the ATE per difference dimension and both counterfactual types per ARG. To recap, ’Aspect’ (Figure 3) measures the number of soundness-related statements, ’Sentiment’ (Figure 4) the density of positive comments, and ’Score’ (Figure 5) the overall rating; all of them compare the ARG’s reviews for the counterfactuals with the one for the original (see Section˜4.3). For an ideal ARG, the ATE for soundness-critical and -neutral edits should lie far apart on all dimensions; the ATE of critical edits should be positive for aspects (soundness-related aspects are added) and negative for sentiment (less positive comments) and score (lower rating).
@@
Although none of the ARGs are significantly affected by flawed research logic, we rank the models to identify tendencies. A coherent ARG should show low ATE variance for soundness-neutral counterfactuals, being resilient to surface-level changes, and a large difference between the ATE for neutral and critical counterfactuals. To capture this in a single score, we express the ATE difference between both conditions as a multiple of the effect standard deviation of the neutral counterfactuals. For each ARG, we compute $z=\frac{|\text{ATE}_{\text{CR}}-\text{ATE}_{\text{NE}}|}{\sigma(\text{ATE}_{\text{NE}})}$ per dimension and then average over all dimensions. This yields the ranking shown in Table˜4. Notably, generic zero-shot prompts perform better than those with guidance, which may act as a distractor. For Phi-4, with its smaller context window, Zero-Guide likely causes truncation of key paper content. While the fine-tuned Reviewer2 ranks highest, DeepReviewer ranks lowest; likely, their backbone models, Llama-2 for Reviewer2 and Phi-4 for DeepReviewer, play an important role since the Zero-Guide results for Phi-4 suggest issues with context size and distractors. This warrants further investigation in future work. Finally, there is no clear link between model size and performance, as Phi-4 and GPT-4.1 both rank highly.
@@
Notably, all ARGs show high standard deviations for soundness-neutral counterfactuals, suggesting that ARGs are sensitive to spurious features similar to issues of prompt sensitivity in LLMs Sclar et al. (2024). To examine this further, we measure lexical similarity using ROUGE-2 and use GPT-4o-mini to detect whether assertions align between the original and soundness-neutral reviews, computing the Jaccard index on assertions. We perform human validation of the automatic alignment with two annotators on $50$ assertion pairs with substantial agreement (Krippendorff’s $\alpha=0.74$) attesting an accuracy of $0.78$. We analyze $48$ randomly sampled papers (roughly 8 per conference). Table 6 in the Appendix reports the detailed results. Surface similarity ranges from $0.25$ ROUGE-2 for DeepReviewer to $0.54$ for Zero-Guide-GPT4om, indicating substantial changes in wording for many ARGs. This pattern holds for the assertion overlap, with a Jaccard index from $0.27$ for DeepReviewer to $0.57$ for Zero-Generic-GPT4om indicating low to moderate overlap in the contents of the reviews. Overall, GPT-4o-mini produces more similar reviews than other ARGs, regardless of the prompting strategy. Even allowing for some measurement error, ARGs remain highly sensitive to spurious surface changes in the input text. Our findings point to another recommendation for future work: all ARG evaluations, whether using human reference data or sensitivity analysis, should report mean performance on multiple soundness-neutral versions of the papers akin to prompt-sensitivity reporting to better reflect their true capability independent of spurious features. Here, our counterfactual dataset may also help improve model consistency through training data augmentation Wu et al. (2021).